OpenStudy (anonymous):

The gas tank of a truck is a cylinder 3.5 ft long with a diameter of 1.5 ft . At the gas station, a pump pours gas at the rate of 5 ft^3 per minute. How many minutes does it take to fill the empty tank with that pump?

OpenStudy (danjs):

First find the volume of the cylinder tank

OpenStudy (danjs):

then divide that by 5 ft^3 / min

OpenStudy (danjs):

\[V = \pi *r^2h\]

OpenStudy (danjs):

They give you the diameter, you need half of that for the radius

OpenStudy (danjs):

radius of the tank = 1.5/2 = 0.75 ft

OpenStudy (danjs):

\[V = \pi *0.75^2*3.5 \approx 6.19~~ft^3\]

OpenStudy (danjs):

The volume of the tank is 6.19 ft^3, and the pump puts out 5 ft^2/min \[minuts = \frac{ 6.19 ~ft^3 }{ 1 }* \frac{ 1~\min }{ 5~ft^3 } \approx 1.24~minutes\]